An unknown solution was analyzed for Ni by an EDTA titration. A 50.00 mL sample of the unknown was treated with 25.00 mL of 0.2404 M EDTA. The excess EDTA was then back titrated with 8.52 mL of 0.0694 M Zn2+ to reach the equivalence point. What was the concentration of Ni (in unit of M) in the 50.00 mL sample? Please keep your answer to three decimal places.
Answer #1

The added EDTA moles and the required Zn + 2 moles are calculated:

n EDTA = M * V = 0.2404 M * 0.025 L = 6.01E-3 mol

n Zn + 2 = 0.0694 M * 0.00852 L = 5.91E-4 mol

The moles of Ni are calculated by difference from the previous moles:

n Ni = 6.01E-3 - 5.91E-4 = 5.42E-3 mol

The molarity is calculated:

[Ni] = n / V = ​​5.42E-3 mol / 0.05 L = 0.108 M

If you liked the answer, please rate it positively, you would help me a lot, thanks.